OpenStudy (anonymous):

State the domain and range of y=-3/4x+4. Please help!

OpenStudy (anonymous):

is the equation: 1) \(\large y=-\frac{3}{4}x +4\) 2) \(\large y=\frac{-3}{4x+4}\) 3) \(\large y=\frac{-3}{4x}+4\) 4) None of the above

OpenStudy (anonymous):

it's 1

OpenStudy (anonymous):

sorry I'm new to the site

OpenStudy (anonymous):

pleeease help!

OpenStudy (anonymous):

Sorry, I was doing dishes.. For the first one, there is no restriction on the values you can use for x.

OpenStudy (anonymous):

So the domain is \((-\infty, \infty)\) or all real numbers.

OpenStudy (anonymous):

The function itself is a line that is descending as x increases. So eventually all numbers will be in the range. So the domain and range are the same.

OpenStudy (anonymous):

so is the answer "all real numbers"? how do I know if it isn't all real numbers? thanks so much by the way.